OpenStudy (anonymous):

If 48 is separated into two parts, and one of the parts is y, what is the other part? i just need to algebraically represent or write an eqaution for.

OpenStudy (anonymous):

It could be any number of things. Do you mean split in two even parts? In which case that would be y=24.

OpenStudy (anonymous):

really im not sure thats all it says is to write an equation for or represent algebraically...

OpenStudy (anonymous):

Well there are a handful of ways you could represent it. you could do like x+y=48 x=-y+48 or y=-x+48

OpenStudy (anonymous):

sure thing thanks.